OpenStudy (anonymous):

\(\large (X)_{i=1}^{\infty} : X _{1} = 1 , X _{2} = 1 ~ , X _{n+1} = X _{n} + 2X _{n-1} \hspace{15px} n \ge 2\) \( \large (Y)_{i=1}^{\infty} : Y _{1} = 1 , Y _{2} = 7 ~ , Y _{n+1} = 2Y _{n} + 3Y _{n-1} \hspace{15px} n \ge 2 \) Prove that none of the terms in those sequences is common (except the first one which is 1)

OpenStudy (anonymous):

ok so X and Y are sequence right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

any by Prove that none of the members in those sequences is common (except the first one which is 1) u mean show that \(X_j\neq Yi ~~~~~~\forall i,j\in N\)

OpenStudy (anonymous):

exactly,and i have got some ideas (i think), write the members , u will discover sth nice about the sequence X ... but for Y i havn't got any...

OpenStudy (anonymous):

ok , i'll see what i got :D

ganeshie8 (ganeshie8):

interesting xD so you want to show that the both sequences have no common terms except 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the funny thing i have found is that \( X _{n} = 2X _{n-1} + (-1)^{n-1} \)

OpenStudy (anonymous):

i'm looking for the same thing for the second seuqence

OpenStudy (anonymous):

yep thats jacobsthal sequence

OpenStudy (anonymous):

ermm

OpenStudy (anonymous):

we can find \(X_{n} \) and \( Y_{n} \) by a way which is called generating function: http://en.wikipedia.org/wiki/Generating_function but i don't think it helps...

OpenStudy (anonymous):

i don't know anything about that..

OpenStudy (anonymous):

We have \[Y_n > X_n\text { for } n > 1\]

OpenStudy (anonymous):

It is easy to prove this by induction. I know we want more than that.

OpenStudy (anonymous):

I will think about it

OpenStudy (anonymous):

but \(X_6>Y_2\) so hmm

OpenStudy (anonymous):

i think @eliassaab mean same members,for example \(X_{2} \) and \( Y_{2} \) or sth else

OpenStudy (anonymous):

yeah i know misunderstanding of the question

OpenStudy (anonymous):

according to http://mathworld.wolfram.com/JacobsthalNumber.html

OpenStudy (anonymous):

\(\Huge X_n=\sum _{r=0}^{\left \lfloor (n-1)/2 \right \rfloor} \binom{n-1-r}{r}2^r \)

OpenStudy (anonymous):

and for \(Y \) ?i don't think it does help...

OpenStudy (anonymous):

so we reduced X_n so far to this \(\large X_n=\frac{1}{3}[2^n-(-1)^n]~~~~~~\forall n\ge 2\)

OpenStudy (anonymous):

think about it i'll be online soon.

OpenStudy (anonymous):

k :D btw u have an answer for this ?

OpenStudy (anonymous):

the interesting thing i note about Y that \(Y_n \mod 10 =1,7,7,5 ~~~~~\text{(in pattern )}\)

OpenStudy (anonymous):

and \(X_n \mod 10 =1,1,3,5 ~~~~~\text{(in pattern )}\)

OpenStudy (anonymous):

so lets assume \(n=4k+q\) these cases that we might have common terms \(n=4k\\n=4k+1\)

OpenStudy (anonymous):

so for n=4k lets assume we have some g,h such that \(X_g=Y_h\) for some \(g=4i\) and \(h=4j\)

OpenStudy (anonymous):

\(\large X_{g}=\frac{1}{3}[2^g-1]\\ ~~~~~~~~\large =\frac{1}{3}[2^{4i}-1]\\ ~~~~~~~~\large =5 \mod 10\\ \large Y_h=2Y_{h-1}+3Y_{h-2} \\ \large ~~~~~=2Y_{4j-1}+3Y_{4j-2}\\ \large ~~~~~ =2\times 7+3 \times 7(\mod 10)=5 \mod 10\)

OpenStudy (anonymous):

ermm we only need a contradiction xD

OpenStudy (anonymous):

ok i give up :P

OpenStudy (michele_laino):

Let's suppose by contradiction that exists an integer, say\[n _{0}>2\] such that we have: \[a _{n _{0}}=b _{n _{0}}\] then we can write: \[\frac{ b_{n _{0}+1}-b _{n _{0}-1} }{ 2 }=a _{n _{0}}+b _{n _{0}-1}\] and: \[a _{n _{0}+1}-a _{n _{0}-1}=a _{n _{0}}+a _{n _{0}-1}\] so if we subtract side by side those above equation, we will get this: \[\frac{ b _{n _{0}+1}-b _{n _{0}-1} }{ 2 }-(a _{n _{n}+1}-a _{n _{0}-1})=b _{n _{0}-1}-a _{n _{0}-1}\] and after some simplification, I get this: \[b _{n _{0}}=a _{n _{0}+1}\] which is a contraddiction!

OpenStudy (michele_laino):

oops..I use a in place of X and b in place of Y

OpenStudy (anonymous):

hmm not sure , note that n need not to be the same for both terms other wise using induction as elissab said works .

OpenStudy (michele_laino):

Please, note that, setting a=b for n_0, I denied the thesis of the statement, and by my subsequent reasoning, I got an absurd statement

OpenStudy (michele_laino):

...so the statement of our question is true!

OpenStudy (anonymous):

@Marki , ur idea was cool ;D , made sense to me ;)

OpenStudy (anonymous):

i'm eager to see ur finished proof.

OpenStudy (anonymous):

i can get the correct proof if u want...these questions are in my number theory book and i can get the answers for all of them (but i don't get and prefer to prove it my self)...

OpenStudy (loser66):

To me, generating function of X is \(X_n= \dfrac{1}{3}(2^n+(-1)^n)\) and generating function of Y is \(Y_n= \dfrac{1}{4}(3^n+(-1)^n)\)

OpenStudy (loser66):

Suppose there is at some value, \(X_n = Y_n\) and we get contradiction. But I didn't get that result yet :)

OpenStudy (michele_laino):

please, note that my solution is logically very consistent!

OpenStudy (loser66):

:) yes, sir @Michele_Laino

OpenStudy (anonymous):

@Loser66 , yes @Marki tried to use generating function but no need to use that! in my book this question is supposed to be solved only bu modular properties.

OpenStudy (michele_laino):

Thank you very much! @Loser66

OpenStudy (anonymous):

@loser OMG ! thank u for the Y function xD i was dying to know it !!!

OpenStudy (anonymous):

@Marki , isn't it \( (-1)^i \) and j?

OpenStudy (anonymous):

\(\large \text{assume there is i,j s.t}\\ \\ \large \begin{align*} X_i& =Y_j~~~~~\text{for some }i,j\in N \\ \frac{1}{3}2^i+(-1)^i &= \frac{1}{4}3^j+(-1)^j \\ 4(2^i+(-1)^i )&=3(3^j+(-1)^j) \\ (2^{i+2}+4(-1)^i ) &= (3^{j+1}+3(-1)^j) \end{align*}\)

OpenStudy (anonymous):

now last line does not give integer solutions i'll show why

OpenStudy (anonymous):

@Marki , ur going very well,continue ;)

OpenStudy (anonymous):

guys wait!!! a very very easy solution!

OpenStudy (anonymous):

we havn't noticed that!

OpenStudy (anonymous):

\( \large (2^{i+2}-3^{j+1} )= 3(-1)^j -4(-1)^i \\\large \text {we have 4 cases ;}\\\large {\color{Red} { \text{ case 1; i is odd j is odd}}}\\ \large (2^{i+2}-3^{j+1} ) =1 \large \\ {\color{Red} {\text{case 2; i is odd j is even}}}\\ \large (2^{i+2}-3^{j+1} ) =-7 \large \\ {\color{Red} {\text{case 3; i is even j is odd}}}\\ \large (2^{i+2}-3^{j+1} ) =7 \large \large \\ {\color{Red} {\text{case 1; i is even j is even}}}\\ (2^{i+2}-3^{j+1} ) =-1 \)

OpenStudy (anonymous):

tell whats ur solution

OpenStudy (anonymous):

see some members of X sequence: 1,1,3,5,11,21,43,85,171,... and if u notice you will discover that members are congruent to 1,1,3,5,3,3,3,5,3,3,.. (mod 8) and here's some members of Y: 1,7,17,55,161,487,1357,... and for (mod 8) we have 1,7,1,7,1,7,.... so none of them can be the same except the first one = 1!

OpenStudy (anonymous):

i see nice !

OpenStudy (anonymous):

thanks,ur work on that functions are good ;)

OpenStudy (anonymous):

its frustrating :P

OpenStudy (anonymous):

well,if u want to solve another question to be calm go to my last question,i added another limit...just think about it tomorrow i'll ask it as a new question,i think it's easy but i havn't got time to work on it.

OpenStudy (anonymous):

well i wont be online :( preparing for final i wish first of feb to prepare some question like these :D

OpenStudy (anonymous):

XD well,goodnight ;) i'm tired and prefer to go to sleep :D

OpenStudy (anonymous):

gn :) sleep well

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@Loser66 ... the sequence (Y) would be: \( Y_{n}=2(3^{n−1})+(−1)^n \) and i checked the answer,for Y4=55 but ur function will give wrong answer ...

ganeshie8 (ganeshie8):

@PFEH.1999 that looks awesome ! you made it really simple by just using the mod properties xD

OpenStudy (anonymous):

@ganeshie8 , thanks...whenever i see this sort of problems this idea come to my mind that try mod 4,8,9,13,17.... :D

OpenStudy (anonymous):

i like 1999 number though :P remineds me of my childhood

OpenStudy (anonymous):

so guys i'll close the question ... thanks @Marki , ur idea was good,it can be used for other problem ;)